OpenStudy (anonymous):

write the expression 256^3/4 in radical notation

OpenStudy (campbell_st):

256 = 2^8 so \[256^{\frac{3}{4}} = (2^8)^{\frac{3}{4}}\] now apply the index law for power of a power.

OpenStudy (campbell_st):

or in radical notation its \[\sqrt[4]{256^3}\]

OpenStudy (campbell_st):

just find the cube of 256

OpenStudy (campbell_st):

I don't know why you need it given its simply 2^6 or 64 when simplified